OpenStudy (anonymous):

Solve 2(3x – 4) = 3x + 7. You must show all work.

OpenStudy (anonymous):

2(3x – 4) = 3x + 7 6x-8=3x+7 DOES THAT HELP?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2(3x-4) = 3x +7 6x - 8 = 3x + 7 6x - 8 - 3x = 7 6x -8 - 3x - 7 = 0 6x - 3x - 8 - 7 = 0 3x - 15 = 0 3x = 15 x = 15 / 3 x = 5